# 1]organization=Center for Research and Formation in Artificial Intelligence, Universidad de los Andes, Colombia, addressline=Carrera 1 No. 18a-12, city=Bogotá, postcode=111711, country=Colombia

### v1

@@
Daniela Ruiz da.ruizl1@uniandes.edu.co    Paula Cárdenas p.cardenasg@uniandes.edu.co    Leonardo Manrique dl.manrique@uniandes.edu.co    Daniela Vega d.vegaa@uniandes.edu.co    Gabriel Mejia gm.mejia@uniandes.edu.co    Pablo Arbeláez pa.arbelaez@uniandes.edu.co
@@ 1 Introduction
Spatial Transcriptomics (ST) is an emerging technology that precisely localizes gene expression profiles within histological images Jiang et al. (2023). While histology analysis is the gold standard for the diagnosis of many diseases Xie et al. (2023), transcriptomics unlocks molecular insights that unveil causal pathways behind pathologies Zeng et al. (2022); Jiang et al. (2023). Beyond disease research, ST has broad applications in developmental biology, enabling the study of tissue formation, cellular differentiation, and organogenesis with spatial resolution Choe et al. (2023). Additionally, ST is valuable in regenerative medicine and tissue engineering, guiding the design of biomaterials and cell-based therapies through a deeper understanding of gene expression patterns in healthy and regenerating tissue Lammi and Qu (2024). By integrating histology with transcriptomics, ST opens a new spectrum of possibilities to understand tissue structure and mechanistic insights into various biological processes Wang et al. (2023).
@@
As with any emerging technology, multiple variations of ST are currently available and under continuous development Stickels et al. (2021); Chen et al. (2015); Ståhl et al. (2016). Notably, as demonstrated by the number of entries in the comprehensive ST repository Wang et al. (2023), Visium Ståhl et al. (2016) has emerged as the most widely used ST technology. The workflow of this technology is depicted in Figure 1 and begins with the preparation of the tissue, where the sample is embedded, sectioned, and placed on a slide with designated capture areas. Next, staining and imaging are performed using standard histological techniques to visualize tissue structures. Once imaged, the tissue is permeabilized, allowing mRNA to be released. Then, this mRNA is captured using barcoded oligonucleotides, enabling spatial mapping of gene expression. A reverse transcription reaction is then used to synthesize cDNA from the captured mRNA, which is subsequently processed into sequencing libraries. Finally, specialized analysis software processes the sequencing data, generating spatially resolved gene expression maps for visualization and interpretation Ståhl et al. (2016).
@@
Despite its advantages, this approach presents key challenges: high costs, the need for domain expertise, and slow adoption in clinical settings, limiting its accessibility in routine diagnostics Pang et al. (2021). In addition to these challenges, on the technical side, it inherits data capturing issues from bulk and single-cell transcriptomics Pham et al. (2023); Avşar and Pir (2023). This problem is known as dropout and corresponds to the failure to detect transcripts even though they are present in the source tissue. In practice, this phenomenon appears as pepper noise in gene expression maps, often requiring single-cell reference datasets to compensate for missing data Avşar and Pir (2023).
@@
Acknowledging these challenges, the deep learning community has delved into democratizing ST by studying gene expression prediction from histology images Jiang et al. (2023). By bypassing the need for specialized sequencing, these approaches offer a more accessible and scalable alternative, enabling patients to obtain molecular insights of a tissue from a standard biopsy image. Leveraging the abundance of public Visium data, multiple deep learning models have emerged to tackle this task He et al. (2020); Pang et al. (2021); Yang et al. (2023, 2024); Xie et al. (2023); Zeng et al. (2022); Mejia et al. (2023). Although these methods consistently report favorable results against the latest state of the art, differences in datasets, preprocessing strategies, and training hyperparameters hinder fair comparisons and compromise the validity of new findings.
Considering the limitations outlined above, we present three key contributions.
@@
    To address the dropout problem, we introduce Spatial transcriptomics Completion with Knowledge from the Local Environment (SpaCKLE), a transformer-based model inspired by the unrivaled power of self-attention mechanisms for next token prediction in natural language processing Dosovitskiy et al. (2020). Notably, SpaCKLE surpasses existing gene completion approaches, achieving a relative 82.5% MSE reduction compared to the median method.
@@ 2.1 Integrated Databases
Recent advances in ST have led to the development of multiple databases. For instance, CROST Wang et al. (2023) is a comprehensive repository with 1033 ST samples from 8 species, 35 tissues, and 56 diseases. Other databases include SpatialDB, Aquila, SPASCER, SODB, and STomicsDB Wang et al. (2023), each offering unique datasets and analytical tools. Although these databases facilitate advanced spatial analyses, they are not specifically designed for the expression profile prediction task.
With the increasing publication of individual datasets obtained with ST technology, database integration has become increasingly important. Similar to the data integration efforts that we present in this work, HEST-1k Jaume et al. (2024) is a recent database that compiles over 1,229 spatial transcriptomic profiles paired with H&E-stained whole-slide images (WSIs). Like SpaRED, HEST-1k provides essential metadata such as organ type and species, and it also extends this information by including disease status, treatment information for cancer samples, and details on the ST technology used. These additional metadata fields facilitate structured analyses across diverse biological conditions. While HEST-1k also presents a python library that provides the possibility of performing data processing, they report not including batch effect correction during their inherent data preprocessing, which may become a limitation for the downstream use of the ST data. Similarly, the newly introduced STimage-1K4M Chen et al. (2024) database contains a diverse collection of 1,149 ST slides, encompassing 4,293,195 spots. Alongside this data, it includes pathologist annotations for 71 slides, providing a valuable resource for assessing clustering methods and dimensionality reduction techniques. Nonetheless, STimage-1K4M lacks any data processing, potentially hindering its use in certain applications.
@@ 2.2 Completion strategies
To address the missing value problem, Mejia et al. (2023) employs a modified adaptive median filter as a completion strategy, replacing dropout values in each gene with the median of a circular region. If multiple spots within the region lack a gene’s value, the method instead uses the median of that gene across the entire WSI. Alternatively, stLearn Pham et al. (2023) uses genetic and morphological similarity to adjust existing spots or predict gene expression for missing values. Moreover, although there are alternative methods that integrate knowledge from single cell RNAseq (Seurat Stuart et al. (2019), Harmony Korsunsky et al. (2019), LIGER Welch et al. (2019) and Tangram Biancalani et al. (2021)), these strategies require a paired single-cell dataset that hampers their usability and practicality. In contrast, SpaCKLE is a reference-free completion method, which stands out from alternatives by leveraging the complete genetic profile of adjacent spots and taking advantage of the transformer capacity to predict missing values.
@@ 2.3 Gene Expression Prediction Benchmarks
A recent study by Jiang et al. (2023) reviews six deep learning methods for gene expression profile prediction, testing their performance on three distinct breast cancer datasets. Although the study provides a solid performance analysis, it focuses exclusively on human breast cancer tissue. Moreover, HEST-1k (Jaume et al. (2024)) also presents a benchmark for the gene expression prediction task. However, instead of evaluating models explicitly designed for gene expression prediction, they assess the effectiveness of histology foundation models in visual feature extraction. Their approach involves using embeddings from 11 different histology foundation models as input to train regression models that predict gene expression. These regression models, based on standard machine learning techniques such as XGBoost, are trained to map histology image embeddings to the log1p-normalized expression levels of preselected genes. Specifically, HEST-1k focuses on predicting the expression of the 50 most highly variable genes across only nine datasets, all derived from cancer patients, to determine the ability of the foundation models to provide the most informative embeddings for gene expression prediction. Additionally, they evaluate model performance exclusively using the Pearson correlation coefficient (PCC), which primarily measures linear associations and may not fully capture how well predictions approximate the actual gene expression values.
Considering the benchmark strategy and focus of both Jiang et al. (2023) and Jaume et al. (2024), we find key differences with the SpaRED benchmark we present in this work. For instance, our benchmark represents a substantial advancement, incorporating an 8.6-fold and 2.3-fold increase in datasets compared to each of these benchmarks. Besides, SpaRED benchmark covers nine different tissue types from both human and mouse subjects with healthy and pathological cases. Figure 2 provides detailed statistics on the number of datasets and spots for each tissue and organism, allowing for a comprehensive evaluation of model generalizability. Additionally, instead of indirectly evaluating histology foundation models through feature extraction, SpaRED directly assesses the performance of state-of-the-art models specifically designed for gene expression prediction. Finally, we include the Mean Square Error (MSE) as an additional performance metric to PCC, to provide a more detailed assessment of each method’s predictive accuracy.
@@ 3.1 Original Datasets and Curation
To build SpaRED, we collect raw data from 7 independent publications Abalo et al. (2021), Parigi et al. (2022), Villacampa et al. (2021), Vicari et al. (2023), Mirzazadeh et al. (2023), Erickson et al. (2022), Fan et al. (2023) and complement them using 5 demonstration datasets from 10X Genomics (available through the SquidPy python package Palla et al. (2022)). We only include datasets with more than one WSI and split the publications’ data by tissue type, resulting in 26 distinct datasets: 14 from human and 12 from mouse, showcasing a variety of tissue samples, as illustrated in Fig. 6.a.
@@
Building on this, we define two types of generalization tasks based on the number of patients in each dataset. Intra-patient generalization considers WSIs as consecutive sections from a single tissue, whereas inter-patient generalization involves WSIs from the same tissue type but across different subjects. To ensure balanced visual distributions, we manually assign WSIs to train, validation, and test sets. In 11 out of 26 cases, a separate test set is defined. For datasets with a limited number of patients or slides, we instead split the data into training and validation sets.
@@
Subsequently, we implement a structured preprocessing pipeline composed of two main stages: data filtering and data processing. In the filtering stage, we begin by setting minimum and maximum cell count thresholds $[10,1,000,000]$, excluding observations that fall outside this range. We then calculate both per-slide and global expression fractions for each gene, retaining only those that meet the minimum expression threshold across spots and the entire dataset. Additionally, we filter out genes with counts outside the predefined range $[10,1,000,000]$ and remove cells with zero expression across all genes.
@@
Following data filtering, we proceed with dataset processing. We normalize gene expression using Transcripts Per Million (TPM) and apply a $log_{2}(x+1)$ transformation. We compute Moran’s I for each gene on each slide to assess spatial autocorrelation, averaging the values across all slides. Based on these scores, we retain the top 128 or 32 genes with the highest spatial autocorrelation, depending on dataset quality. Finally, we apply ComBat batch correction to mitigate batch effects.
@@
As a result, the final SpaRED dataset includes 105 slides and 308,843 spots from 35 patients. Table 1 provides a comprehensive breakdown of the dataset statistics, which also presents the proportion of missing data before and after processing. The table illustrates how the proposed processing steps effectively clean the data, substantially reducing missing values.
@@
Table 1: Detailed overview of the SpaRED datasets, showcasing the generalization task, the number of genes analyzed, the abbreviation for each dataset, the organism studied, the tissue or disease, and the number of slides, patients and spots. Additionally, the table presents key statistics on data corruption and missing values. The ‘Corrupt Spots’ column shows the percentage of spots with at least one corrupted value in each dataset. ‘Missing data before’ indicates the proportion of missing data prior to any processing, while ‘Missing data after’ reflects the percentage remaining after processing.
@@
Slides Patients Spots
@@
Inter- Patient 128 VMB [20] Mouse Brain 14 4 43804 100% 89% 28%

### v2

@@
Daniela Ruiz da.ruizl1@uniandes.edu.co    Paula Cárdenas p.cardenasg@uniandes.edu.co    Leonardo Manrique dl.manrique@uniandes.edu.co    Daniela Vega d.vegaa@uniandes.edu.co    Gabriel M. Mejia gm.mejia@uniandes.edu.co    Pablo Arbeláez pa.arbelaez@uniandes.edu.co
@@ 1 Introduction
Spatial Transcriptomics (ST) is an emerging technology that precisely localizes gene expression profiles within histological images (Jiang et al., 2023). While histology analysis is the gold standard for the diagnosis of many diseases (Xie et al., 2023), transcriptomics unlocks molecular insights that unveil causal pathways behind pathologies (Zeng et al., 2022; Jiang et al., 2023). Beyond disease research, ST has broad applications in developmental biology, enabling the study of tissue formation, cellular differentiation, and organogenesis with spatial resolution (Choe et al., 2023). Additionally, ST is valuable in regenerative medicine and tissue engineering, guiding the design of biomaterials and cell-based therapies through a deeper understanding of gene expression patterns in healthy and regenerating tissue (Lammi and Qu, 2024). By integrating histology with transcriptomics, ST opens a new spectrum of possibilities to understand tissue structure and mechanistic insights into various biological processes (Wang et al., 2023).
@@
As with any emerging technology, multiple variations of ST are currently available and under continuous development (Stickels et al., 2021; Chen et al., 2015; Ståhl et al., 2016). Notably, as demonstrated by the number of entries in the comprehensive ST repository (Wang et al., 2023), Visium (Ståhl et al., 2016) has emerged as the most widely used ST technology. The workflow of this technology is depicted in Fig. 1 and begins with the preparation of the tissue, where the sample is embedded, sectioned, and placed on a slide with designated capture areas. Next, staining and imaging are performed using standard histological techniques to visualize tissue structures. Once imaged, the tissue is permeabilized, allowing mRNA to be released. Then, this mRNA is captured using barcoded oligonucleotides, enabling spatial mapping of gene expression. A reverse transcription reaction is then used to synthesize cDNA from the captured mRNA, which is subsequently processed into sequencing libraries. Finally, specialized analysis software processes the sequencing data, generating spatially resolved gene expression maps for visualization and interpretation (Ståhl et al., 2016).
@@
Despite its advantages, this approach presents key challenges: high costs, the need for domain expertise, and slow adoption in clinical settings, limiting its accessibility in routine diagnostics (Pang et al., 2021). In addition to these challenges, on the technical side, it inherits data capturing issues from bulk and single-cell transcriptomics (Pham et al., 2023; Avşar and Pir, 2023). This problem is known as dropout and corresponds to the failure to detect transcripts even though they are present in the source tissue. In practice, this phenomenon appears as pepper noise in gene expression maps, often requiring single-cell reference datasets to compensate for missing data (Avşar and Pir, 2023).
@@
Acknowledging these challenges, the deep learning community has delved into democratizing ST by studying gene expression prediction from histology images (Jiang et al., 2023). By bypassing the need for specialized sequencing, these approaches offer a more accessible and scalable alternative, enabling subjects to obtain molecular insights of a tissue from a standard biopsy image. Leveraging the abundance of public Visium data, multiple deep learning models have emerged to tackle this task (He et al., 2020; Pang et al., 2021; Yang et al., 2023, 2024; Xie et al., 2023; Zeng et al., 2022; Mejia et al., 2023). Although these methods consistently report favorable results against the latest state of the art, differences in datasets, preprocessing strategies, and training hyperparameters hinder fair comparisons and compromise the validity of new findings.
In our previous MICCAI paper titled "Enhancing Gene Expression Prediction from Histology Images with Spatial Transcriptomics Completion" (Mejia et al., 2024), we introduced initial efforts to address the limitations discussed above. In this work, we substantially build upon and refine those initial contributions. First, we enhance the methodology by introducing comprehensive ablation studies to support our design choices for SpaCKLE, including the contribution of data pre-completion, the integration of visual features, the effect of incorporating context genes information, and the impact of neighborhood size. Second, we broaden the SpaRED benchmark by adding the state-of-the-art model HGGEP (Li et al., 2024) and systematically evaluating its performance across all 26 datasets. Third, we provide a more comprehensive analysis with additional qualitative and statistical results for both our completion model and the SpaRED Benchmark, offering more profound insights into SpaCKLE’s performance and a more detailed comparative evaluation of existing gene expression prediction models.
Our key contributions can be summarized as follows.
@@
    To address the dropout problem, we introduce Spatial transcriptomics Completion with Knowledge from the Local Environment (SpaCKLE), a transformer-based model inspired by the unrivaled power of self-attention mechanisms for next token prediction in natural language processing (Dosovitskiy et al., 2020). Notably, SpaCKLE surpasses existing gene completion approaches, achieving a relative 82.5% MSE reduction compared to the median method.
@@ 2.1 Integrated Databases
Recent advances in ST have led to the development of multiple databases. For instance, CROST (Wang et al., 2023) is a comprehensive repository with 1033 ST samples from 8 species, 35 tissues, and 56 diseases. Other databases include SpatialDB, Aquila, SPASCER, SODB, and STomicsDB (Wang et al., 2023), each offering unique datasets and analytical tools. Although these databases facilitate advanced spatial analyses, they are not specifically designed for the expression profile prediction task.
With the increasing publication of individual datasets obtained with ST technology, database integration has become increasingly important. Similar to the data integration efforts that we present in this work, HEST-1k (Jaume et al., 2024) is a recent database that compiles over 1,229 spatial transcriptomic profiles paired with H&E-stained whole-slide images (WSIs). Like SpaRED, HEST-1k provides essential metadata such as organ type and species, and it also extends this information by including disease status, treatment information for cancer samples, and details on the ST technology used. These additional metadata fields facilitate structured analyses across diverse biological conditions. While HEST-1k also presents a python library that provides the possibility of performing data processing, they report not including batch effect correction during their inherent data preprocessing, which may become a limitation for the downstream use of the ST data. Similarly, the newly introduced STimage-1K4M (Chen et al., 2024a) database contains a diverse collection of 1,149 ST slides, encompassing 4,293,195 spots. Alongside this data, it includes pathologist annotations for 71 slides, providing a valuable resource for assessing clustering methods and dimensionality reduction techniques. Nonetheless, STimage-1K4M lacks any data processing, potentially hindering its use in certain applications.
@@ 2.2 Completion strategies
Several strategies have been proposed to address missing data in spatial transcriptomics, broadly falling into two categories: reference-based and reference-free methods.
Reference-based methods integrate spatial transcriptomics with matching single-cell RNA-seq datasets to infer missing gene expression and enhance resolution. For instance, Tangram (Biancalani et al., 2021) aligns spatial and single-cell transcriptomic profiles to map gene expression from the single-cell domain onto tissue. SpaGE (Abdelaal et al., 2020) projects spatial data into a latent space constructed from single-cell references to predict unmeasured genes. Seurat (Stuart et al., 2019), Harmony (Korsunsky et al., 2019), LIGER (Welch et al., 2019), gimVI (Lopez et al., 2019), and stPlus (Shengquan et al., 2021) follow similar principles, combining multimodal alignment or probabilistic modeling to impute spatial gene expression using external scRNA-seq data.
While effective, these methods require carefully curated, high-quality single-cell datasets from the same tissue and condition—resources that are often difficult to obtain. Moreover, utilizing an scRNA-seq reference, increases the resources needed to clean the ST dataset at the risk of inducing batch effects due to dissimilar sequencing technologies (Marel, 2024). Additionally, reference-based models depend on alignment quality, which remains suboptimal despite ongoing advancements, potentially introducing bias in the completed data (Yan et al., 2024).
Reference-free approaches, in contrast, rely exclusively on the spatial structure and expression context within each ST slide. For example, SEPAL (Mejia et al., 2023) uses a modified adaptive median filter to replace dropout values with the median expression in a local circular region; if the region lacks sufficient data, the method falls back to the global median. Alternatively, stLearn (Pham et al., 2023) uses genetic and morphological similarity to adjust existing spots or predict gene expression for missing values.
SpaCKLE is a reference-free completion method, which stands out from alternatives by leveraging the complete genetic profile of adjacent spots and taking advantage of the transformer capacity to predict missing values. SpaCKLE captures local gene-gene and spot-spot dependencies, allowing it to complete missing values using only intrinsic spatial information. This makes it especially suited for Visium datasets lacking a corresponding single-cell reference, offering greater flexibility and broader applicability.
@@ 2.3 Gene Expression Prediction Benchmarks
A recent study by Jiang et al. (2023) reviews six deep learning methods for gene expression profile prediction, testing their performance on 3 distinct breast cancer datasets. Although the study provides a solid performance analysis, it focuses exclusively on human breast cancer tissue. Moreover, HEST-1k (Jaume et al., 2024) also presents a benchmark for the gene expression prediction task. However, instead of evaluating models explicitly designed for gene expression prediction, they assess the effectiveness of histology foundation models in visual feature extraction. Their approach involves using embeddings from 11 different histology foundation models as input to train regression models that predict gene expression. These regression models, based on standard machine learning techniques such as XGBoost, are trained to map histology image embeddings to the log1p-normalized expression levels of preselected genes. Specifically, HEST-1k focuses on predicting the expression of the 50 most highly variable genes across only nine datasets, all derived from human cancer samples, to determine the ability of the foundation models to provide the most informative embeddings for gene expression prediction. Additionally, they evaluate model performance exclusively using the Pearson correlation coefficient (PCC), which primarily measures linear associations and may not fully capture how well predictions approximate the actual gene expression values.
Considering the benchmark strategy and focus of both Jiang et al. (2023) and Jaume et al. (2024), we find key differences with the SpaRED benchmark we present in this work. While Jaume et al. (2024) includes a greater total number of datasets, slides, and spots, it reports results for only 9 datasets. In contrast, SpaRED reports results on 26 datasets, which is 2.8 times more than Jaume et al. (2024) and 8.6 times more than Jiang et al. (2023). Additionally, SpaRED benchmark covers nine different tissue types from both human and mouse subjects with healthy and pathological cases, in contrast to Jiang et al. (2023). Fig. 2 provides detailed statistics on the number of datasets and spots for each tissue and organism, allowing for a comprehensive evaluation of model generalizability. Additionally, instead of indirectly evaluating histology foundation models through feature extraction, SpaRED directly assesses the performance of state-of-the-art models specifically designed for gene expression prediction. Finally, we include the Mean Square Error (MSE) as an additional performance metric to PCC, to provide a more detailed assessment of each method’s predictive accuracy.
@@ 3.1 Original Datasets and Curation
To build SpaRED, we collect raw data from 7 independent publications (Abalo et al., 2021), (Parigi et al., 2022), (Villacampa et al., 2021), (Vicari et al., 2023), (Mirzazadeh et al., 2023), (Erickson et al., 2022), (Fan et al., 2023) and complement them using 5 demonstration datasets from 10X Genomics (available through the SquidPy python package (Palla et al., 2022)). We only include datasets with more than one WSI and split the publications’ data by tissue type, resulting in 26 distinct datasets: 14 from human and 12 from mouse, showcasing a variety of tissue samples, as illustrated in Fig. 2.
@@
Building on this, we define two types of generalization tasks based on the number of subjects in each dataset. Intra-subject generalization considers WSIs as consecutive sections from a single tissue and subject, whereas inter-subject generalization involves WSIs from the same tissue type but across different subjects. To ensure balanced visual distributions, we manually assign WSIs to train, validation, and test sets. In 11 out of 26 cases, a separate test set is defined. For datasets with a limited number of subjects or slides, we instead split the data into training and validation sets.
@@
Subsequently, we implement a structured preprocessing pipeline composed of two main stages: data filtering and data processing. In the filtering stage, we begin by setting minimum and maximum cell count thresholds $[10,1000000]$, excluding observations that fall outside this range. We then calculate both per-slide and global expression fractions for each gene, retaining only those that meet the minimum expression threshold across spots and the entire dataset. Additionally, we filter out genes with counts outside the predefined range $[10,1000000]$ and remove cells with zero expression across all genes. This initial filtering step is intentionally aggressive, eliminating the vast majority of low-expression or non-informative genes to improve data quality and reduce noise.
@@
Following data filtering, we proceed with dataset processing. We normalize gene expression using Transcripts Per Million (TPM) and apply a $log_{2}(x+1)$ transformation. We compute Moran’s I for each gene on each slide to assess spatial autocorrelation, averaging the values across all slides. Inspection of the full Moran’s I ranking revealed that genes falling at the bottom exhibited markedly higher proportions of missing data. Consequently, we retain the top 128 or 32 genes with the highest spatial autocorrelation. We select these numbers to ensure a consistently high-quality set of informative genes and computational efficiency across all datasets. However, the number of genes is customizable when using the SpaRED library for dataset creation. Finally, we apply ComBat (Johnson et al., 2007) batch correction to mitigate batch effects.
@@
As a result, the final SpaRED dataset includes 105 slides and 308,843 spots from 35 subjects. Table 1 provides a comprehensive breakdown of the dataset statistics. Moreover, it shows the proportion of missing data before and after processing. This proportion refers to the fraction of spatial spots with missing expression values for any of the 32 or 128 retained genes in each dataset. The table 1 demonstrates that our processing pipeline effectively cleans the data, substantially reducing the amount of missing values.
@@
Table 1: Detailed overview of the SpaRED datasets, showcasing the generalization task, the number of genes analyzed, the abbreviation for each dataset, the organism studied, the tissue or disease, and the number of slides, subjects and spots. Additionally, the table presents key statistics on data corruption and missing values. The ‘Corrupt Spots’ column shows the percentage of spots with at least one corrupted value in each dataset. ‘Missing data before’ indicates the proportion of missing data prior to any processing, while ‘Missing data after’ reflects the percentage remaining after processing.
@@
Slides Subjects Spots
@@
Inter-
Subject
128 VMB [20] Mouse Brain 14 4 43804 100% 89% 28%